I don’t understand what I did wrong. Could you please explain why it’s 32W/m^2?
Question 16 The speed of sound in air is 340 m/s, and the density of air is 1.2 kg/m3. If the displacement amplitude of a 440-Hz sound wave is 10 μm, what is the intensity of the wave? 16 W/m2 rrect Answer 32 W/m2 0.32 W/m 12 W/n You Answered :0.16 W/m2 2/2 pts Question 17
Answer #1

2.7 -6 11-27 fascal 、au, ound ware 2一 2 = (11-27 8 16 0.155 6 = 0-16 wrm.